A publisher reports that 33% of their readers own a particular make of car. A marketing executive wants to test the claim that the percentage is actually different from the reported percentage. A random sample of 100 found that 24% of the readers owned a particular make of car. Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.33

Ha : p 0.33

n = 100

= 0.24

P0 = 0.24

1 - P0 = 1 - 0.24 = 0.76

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.33 - 0.24/ [0.24 * 0.76 / 100 ]

= 2.107

Test statistic = z = 2.107

P-value = 0.0351